# Understanding the differences in Foundation Models: Attention, State Space Models, and Recurrent Neural Networks

### v2

@@ 3.2.2 State Space Models
SSM models are straightforward to rewrite in the DSF given their intrinsic recurrent linear representation. However, similarly to attention, we slightly rewrite the standard representation introduced in the literature in order to reveal deeper insights obscured by the standard representation focused on computational efficiency. The detailed derivation can be found in Appendix D.
@@
with $\Delta_{i}=\textrm{diag}(\textrm{softplus}(W_{\Delta}(W_{u}u_{i})+b_{\Delta}))\in\mathbb{R}^{d\times d}$, $b_{i}=W_{B}u_{i}\in\mathbb{R}^{n},\,c_{i}=W_{C}u_{i}\in\mathbb{R}^{n}$, and $W_{u}\in\mathbb{R}^{p\times d},\,W_{\Delta}\in\mathbb{R}^{d\times p}$ are weight matrices with $p<d$, and $b_{\Delta}\in\mathbb{R}^{d}$ is a bias.
Note that in formulation (18) the dimensions of the matrices are $\Lambda_{i}\in\mathbb{R}^{nd\times nd}$, $B_{i}\in\mathbb{R}^{nd\times d}$, $C_{i}\in\mathbb{R}^{d\times nd}$, i.e., $n$ is the state dimension and $d$ is the input dimension in the original formulation (3).
@@ 3.2.3 Recurrent Neural Networks
Given their recurrent nature, one can express LSTMs (5) in the DSF with some basic algebraic manipulations (see Appendix E for details). Once again, we slightly rewrite the standard representation since our goal is to obtain mathematical insights as opposed to computational efficiency.
@@
$B_{i}$ $=\mathrm{diag}(\sigma(W_{i}u_{i}))\odot W_{u}\in\mathbb{R}^{d\times d},$ (19b)
$C_{i}$ $=\mathrm{diag}(\sigma(W_{o}u_{i}))\in\mathbb{R}^{d\times d},$ (19c)
where $W_{f},W_{i},W_{o},W_{u}\in\mathbb{R}^{d\times d}$ are the learnt parameters in (8).
It is important to note that here the dimension of the hidden state $h_{i}$ is equal to the number of input channels $d$. This is in contrast with attention and SSMs, where the dimension of the hidden state $h_{i}$ in the DSF (11) is $nd$. In the case of qLSTMs, $n=1$. This fact will become relevant in further discussions; we refer to the fact that $n>1$ as state expansion.
@@
Given the similarities of the RG-LRU model [De et al., 2024] with a SSM, it is straightforward to reformulate it into the DSF (11) without the need for modifications besides simple algebraic manipulations. Hence, the RG-LRU can be expressed in the DSF as
$\Lambda_{i}$ $=e^{-cr_{i}\odot\mathrm{softplus}(A)}\in\mathbb{R}^{d\times d},$ (20a)
$B_{i}$ $=\sqrt{1-\Lambda_{i}^{2}}\odot\mathrm{diag}(\sigma(W_{B}u_{i}))\in\mathbb{R}^{d\times d},$ (20b)
$C_{i}$ $=\mathbb{I}_{d},$ (20c)
@@ 4 Architecture Comparison: Theoretical and Experimental Results
[…]
@@
Separable attention is used to avoid computation of the query-key matrix $\mathbf{q}\mathbf{k}^{\top}$. It allows to compute $\mathbf{k}^{\top}\mathbf{v}$ before multiplying the queries $\mathbf{q}$, which reduces the computational complexity from quadratic to linear in sequence length. The DSF makes apparent how a necessary condition for separability is for the map $\zeta(\cdot)$ to be expressed by a finite-dimensional kernel. However, in the case of $\texttt{softmax}(\cdot)$, an infinite-dimensional kernel is needed, i.e., in the DSF softmax attention requires $n=\infty$. This insight can mathematically explain why the good performance observed in softmax attention can only be approximated by separable attention mechanisms, SSMs, or RNNs; but no other architecture is equivalent. The DSF predicts that softmax can be better approximated by growing $n$, which we show in the following theoretical result:
@@
The result follows from the fact that the first $N$ states and the output in (11) can be chosen to be independent of the additional states. The proof is given in Appendix H. ∎
@@
Since $N=nd$, this means that the larger the state expansion $n$, the more expressivity the architecture has – in the case of softmax attention (2), $n=\infty$ (Lemma 1). To show expressivity with increasing $n$ empirically, we compare linear attention (16), for which it holds that $n=m$, and softmax attention on the MQAR. Figure 1 shows that with larger state expansion linear attention converges to the performance of softmax attention, which achieves perfect accuracy throughout.
@@
For this comparison, we assume a generalized linear attention with $\phi(q_{i})=q_{i}$, $\psi(k_{j})=k_{j}$, and general normalization function $\eta(q_{i},\mathbf{k}_{i})$. By comparing the DSF expressions for both Generalized Linear Attention (15) and S6 (18), one can immediately see that the S6 parameters $b_{i}=W_{B}u_{i}\in\mathbb{R}^{n},\,c_{i}=W_{C}u_{i}\in\mathbb{R}^{n}$ directly correspond with keys and queries in attention, i.e. $k_{i}=b_{i}$ and $q_{i}=c_{i}$. Moreover, the state expansion $n$ in S6 is the same as the hidden dimension in attention. However, while this leads to an equivalent output matrix $C_{i}$ in the DSF parametrization for both architectures, there are remarkable differences between the two:
  • Number of parameters. The transition matrix $\Lambda_{i}$ has $d$ parameters in S6 (18) and only $1$ in attention. In attention (15), $\eta(q_{i},\mathbf{k})$ is the only parameter in $\Lambda_{i}$, and it is by definition a scalar. In S6, the parameters in $\Lambda_{i}$ are determined by $\Delta_{i}\in\mathbb{R}^{d}$, which has $d$ different parameters. However, it was shown in [Dao and Gu, 2024] that the number of parameters in $\Lambda_{i}$ can be reduced to $1$ – similar to attention – without compromising performance.
@@
We remark that multi-headed attention increases the number of parameters in $\Lambda_{i}$ from $1$ to the number of heads $s$; for more details see Appendix F. While the number of parameters in the state transition $\Lambda_{i}$ does play a role in increasing performance (multi-headed attention typically performs better than single-headed attention [Vaswani et al., 2017]), the results in [Dao and Gu, 2024] suggest that this role is small. The larger influence thus lies in the recursive structure of $\Lambda_{i}$ and $B_{i}$ and/or the parameterization of normalization $\eta(\cdot)$. To further investigate this and the role of normalization in attention, we compare S6 and softmax attention to SSD [Dao and Gu, 2024], linear attention [Katharopoulos et al., 2020], and normalized attention on the multi-query associative recall (MQAR) benchmark [Arora et al., 2023] and train the three attention-based methods on WikiText-103. Inspired by S6, we define normalized attention as the attention function
@@
where $W_{\eta}\in\mathbb{R}^{1\times d}$ is an additional learnt parameter. The MQAR results are shown in Figure 2 and the training perplexity on WikiText-103 in Table 1.
[…]
Table 1: Training perplexity score for different Attention architectures (70M params) on the WikiText-103 corpus.
@@
In the comparison of RNNs and S6 it is immediate to observe several similarities. In particular, as shown in Appendix G, the state transition matrix $\Lambda_{i}$ in S6 (18) can be rewritten when assuming $A=a\cdot\mathbb{I}_{nd}$ as
@@
Notice that when no state expansion is considered, i.e., $n=1$ and $\mathbb{I}_{n}=1$, this expression almost coincides with the qLSTM state transition (19a), with the only difference that (I) it uses the reversed sigmoid function instead of a sigmoid for the forget gate, and (II) there is an additional learnt parameters $a$ in the exponent. As shown in the recent xLSTM paper [Beck et al., 2024], state expanded LSTMs [^3] The presented state expanded LSTM versions, cannot be directly translated into the DSF framework, since the gates not only depend on the inputs $u_{i}$ but also on past outputs $y_{i-1}$. However, the used state expansion is essentially $n=d$, hence leading to a DSF system of size $d^{2}$. can yield similar performance to S6. This aligns with Lemma 2, and further highlights the importance of state expansion for expressivity. In the case of qLSTM and RG-LRU, state expansion can be easily incorporated by changing the dimensions of the projections $W_{f},\,W_{i},\,W_{o}$, where the $\odot$ operation in RG-LRU would be replaced by blockwise operations $\otimes$. Finally, the most apparent difference between the two RNN variants qLSTM and RG-LRU and S6 is the parameter coupling in $\Lambda_{i}$ and $B_{i}$. While qLSTM does not use a coupling, the couplings in RG-LRU and S6 are performed with different nonlinearities, which is discussed in more detail in [De et al., 2024, Appendix A].
Inspired by the subtle difference in the state transition, we compare the original qLSTM state transition and the S6 state transition on the MQAR benchmark. Specifically, we compare qLSTM (8) and qLSTM with only the state transition replaced by (22). The performance of both models is shown in Figure 3. We note that the reversed sigmoid state transition outperforms the original state transition on all three benchmark tasks, i.e., the performance of qLSTMs can be improved by insights from S6.
Figure 3: Comparison of qLSTM (8) and a qLSTM variant where the original state transition $\Lambda_{i}$ is replace by (22).
@@ 5 Related Work
State-space models emerged from the S4 architecture by Gu et al. [2022a], who developed a new theoretically principled approach to sequence modeling rooted in polynomial approximation theory [Gu et al., 2020]. The result is a transformer-like architecture [Vaswani et al., 2017], where attention is replaced by a linear recurrent neural network with special reparametrization. The design of S4 got later simplified in [Gupta et al., 2022, Gu et al., 2022b], who achieve state-of-the-art on the long-range arena (LRA) [Tay et al., 2021] with a highly efficient recurrent mechanism leveraging convolutional views [Li et al., 2022], or parallel scans [Smith et al., 2023, Orvieto et al., 2023].
@@
The high efficiency of SSMs (linear processing) makes them particularly appealing when compared to softmax attention-based transformers, where both inference time and memory suffer quadratically from sequence length. The S4 architecture found first successful applications in reinforcement learning [Lu et al., 2023], vision [Nguyen et al., 2022], audio [Goel et al., 2022] as well as online learning [Zucchet et al., 2023b]. Initial attempts in language modeling [Fu et al., 2023, Wang et al., 2022], supported by theoretical investigations [Orvieto et al., 2024, Wang and Xue, 2023] hint at some necessary architectural improvements to unlock the NLP domain. Leveraging in-context learning arguments, a few works [Peng et al., 2023, Sun et al., 2023, Katsch, 2023] started incorporating input selectivity mechanisms [Olsson et al., 2022] into SSMs. These efforts culminated in the Mamba architecture [Gu and Dao, 2023], which proposed an highly efficient and light (in terms of parameters) input selectivity strategy, with drastic improvements when comparing to earlier variants (H3 [Fu et al., 2023] and Hyena [Poli et al., 2023]) on text. This approach was also shown to be effective at byte level [Wang et al., 2024]. Beyond text, Mamba was recently applied to the vision domain [Liu et al., 2024b, Zhu et al., 2024] – with outstanding results compared to ViTs [Touvron et al., 2022] both in terms of performance and efficiency. Other applications include e.g. genetics [Schiff et al., 2024], and point clouds [Liu et al., 2024a]. Further, improvements on architectural aspects were proposed in [Yang et al., 2023, Dao and Gu, 2024, Qin et al., 2024].
@@
The design of Mamba is also strongly supported by theoretical evidence showing precisely its superior expressive power compared to S4 [Cirone et al., 2024]. This boost in computational power is due to Mamba’s novel input selectivity mechanism resembling gating, which unlocks content-dependent reasoning [Gu and Dao, 2023, Olsson et al., 2022]. Interestingly, input selectivity brings SSMs closer to attention: in particular, Ali et al. [2024] showed that the particular parametrization of Mamba can be linked to a non-normalized softmax operator. This finding is also supported by evidence from language theory – Mamba and Attention can solve a similar class of problems [Merrill et al., 2024]. Beyond Ali et al. [2024] the connection between linear attention and linear RNNs has been illustrated a few times in the literature [Katharopoulos et al., 2020, Schlag et al., 2021, Zucchet et al., 2023a]. Compared to these works and to Ali et al. [2024], we offer here a more careful comparison identifying some precise distinctions between Mamba, linear, and softmax attention – which play a nontrivial role in practice and can help bring to light interesting architectural improvements.

### v3

@@ 3.2.2 State Space Models
SSM models are straightforward to rewrite in the DSF given their intrinsic recurrent linear representation. However, similarly to attention, we slightly rewrite the standard representation introduced in the literature to reveal deeper insights obscured by the standard representation focused on computational efficiency. The detailed derivation can be found in Appendix E.
@@
with $\Delta_{i}=\textrm{diag}(\textrm{softplus}(W_{\Delta}(W_{u}u_{i})+b_{\Delta}))\in\mathbb{R}^{d\times d}$, $b_{i}=W_{B}u_{i}\in\mathbb{R}^{n},\,c_{i}=W_{C}u_{i}\in\mathbb{R}^{n}$, and $W_{u}\in\mathbb{R}^{p\times d},\,W_{\Delta}\in\mathbb{R}^{d\times p}$ are weight matrices with $p<d$, and $b_{\Delta}\in\mathbb{R}^{d}$ is a bias. Note that in formulation (19) the dimensions of the matrices are $\Lambda_{i}\in\mathbb{R}^{nd\times nd}$, $B_{i}\in\mathbb{R}^{nd\times d}$, $C_{i}\in\mathbb{R}^{d\times nd}$, i.e., $n$ is the state dimension and $d$ is the input dimension in the original formulation (4).
@@ 3.2.3 Recurrent Neural Networks
Given their recurrent nature, one can express LSTMs (5) in the DSF with some basic algebraic manipulations (see Appendix F for details). Once again, we slightly rewrite the standard representation since our goal is to obtain mathematical insights as opposed to computational efficiency.
@@
$B_{i}$ $=\mathrm{diag}(\sigma(W_{i}u_{i}))\odot W_{u}\in\mathbb{R}^{d\times d},\quad C_{i}=\mathrm{diag}(\sigma(W_{o}u_{i}))\in\mathbb{R}^{d\times d},$ (20b)
where $W_{f},W_{i},W_{o},W_{u}\in\mathbb{R}^{d\times d}$ are the learnt parameters in (8). It is important to note that here the dimension of the hidden state $h_{i}$ is equal to the number of input channels $d$, whereas in attention and SSMs the dimension of the hidden state $h_{i}$ in the DSF (11) is $nd$. For qLSTMs $n=1$, which will become relevant in further discussions; we refer to the fact that $n>1$ as state expansion.
@@
Given the similarities of RG-LRU [10] and SSMs, it is straightforward to reformulate it into the DSF (11) without the need for modifications besides simple algebraic manipulations. Hence, the RG-LRU can be expressed in the DSF as
$\Lambda_{i}$ $=e^{-cr_{i}\odot\mathrm{softplus}(A)}\in\mathbb{R}^{d\times d},\quad B_{i}=\sqrt{1-\Lambda_{i}^{2}}\odot\mathrm{diag}(\sigma(W_{B}u_{i}))\in\mathbb{R}^{d\times d},\quad C_{i}=\mathbb{I}_{d},$ (21a)
@@ 4 Architecture Comparison: Theoretical and Experimental Results
In this section, we use the DSF to explore some of the long-standing questions between attention, SSMs, and RNNs. We provide theoretical results and/or numerical experiments to substantiate our claims. The experiments presented below are performed on the multi-query associative recall (MQAR) [24] and long range arena (LRA) [3] benchmarks using the code bases [^2] https://github.com/HazyResearch/zoology;  https://github.com/google-research/long-range-arena provided with the benchmarks. The complete experimental setup and computational resources used are detailed in Appendices J and K, respectively, and a statistical analysis is provided in Appendix L.
@@
Separable attention is used to avoid computation of the query-key matrix $\mathbf{q}\mathbf{k}^{\top}$. It allows to compute $\mathbf{k}^{\top}\mathbf{v}$ before multiplying the queries $\mathbf{q}$, which reduces the computational complexity from quadratic to linear in sequence length. While the DSF shows how separable attention, and in particular kernelized attention can be rewritten as a recurrence (11), such a reformulation is only practical for a finite state dimension. However, in the case of $\texttt{softmax}(\cdot)$, an infinite-dimensional kernel is needed, i.e., in the DSF, softmax attention requires $n=\infty$. This insight can mathematically explain why the good performance observed for softmax attention can only be approximated by separable attention mechanisms, SSMs, or RNNs; but no other architecture is equivalent. The DSF predicts that softmax can be better approximated by growing $n$, which we show in the following theoretical result.
@@
The result follows from the fact that the first $N$ states and the output in (11) can be chosen to be independent of the additional states. The full proof is given in Appendix D. ∎
@@
Therefore, it holds that the expressivity of a model is non-decreasing with increasing state expansion $n$ (and state dimension $N=nd$), if the rest of the architecture is held constant. As the softmax attention has an infinite hidden state dimension, i.e. $n=\infty$ (Lemma 1), we investigate empirically how its performance compares to linear attention (16), with increasing state dimension on the MQAR. Figure 1 shows that with larger $n$ linear attention converges to the performance of softmax attention, which achieves perfect accuracy throughout.
[…]
Model LRA WikiText
Linear Att. (16) 53.52 17.42
Norm. Att. (22) 58.08 16.43
Softmax Att. (2) 55.96 13.15
S6 (4) 66.84 N/A
Table 1: Average accuracy on the LRA benchmark and training perplexity score for different attention architectures (70M params) on the WikiText-103 corpus.
@@
By comparing the DSF expressions for both generalized linear attention (15) and S6 (19), we notice that the S6 parameters $b_{i}=W_{B}u_{i}\in\mathbb{R}^{n},\,c_{i}=W_{C}u_{i}\in\mathbb{R}^{n}$ directly correspond to the keys and queries in attention, i.e. $k_{i}=b_{i}$ and $q_{i}=c_{i}$. Moreover, the state expansion $n$ in S6 is the same as the hidden dimension in attention. However, while this leads to an equivalent output matrix $C_{i}$ in the DSF parametrization for both architectures, there are remarkable differences between the two:
  • Number of parameters. The transition matrix $\Lambda_{i}$ has $d$ parameters in S6 (19) and only $1$ in attention. In attention (15), $\eta(q_{i},\mathbf{k})$ is the only parameter in $\Lambda_{i}$, and it is by definition a scalar. In S6, the parameters in $\Lambda_{i}$ are determined by $\Delta_{i}\in\mathbb{R}^{d}$, which has $d$ different parameters. However, it was shown in [8] that the number of parameters in $\Lambda_{i}$ can be reduced to $1$ – similar to attention – without compromising performance. Note that multi-headed attention increases the number of parameters in $\Lambda_{i}$ from $1$ to the number of heads $s$; for more details see Appendix G.
@@
While the number of parameters in the state transition $\Lambda_{i}$ does play a role in increasing performance (multi-headed attention typically performs better than single-headed attention [2]), the results in [8] suggest that this role is small. The larger influence thus lies in the recursive structure of $\Lambda_{i}$ and $B_{i}$ and/or the parameterization of normalization $\eta(\cdot)$. To further investigate this and the role of normalization in attention, we compare S6 and softmax attention to SSD [8], linear attention [14], and normalized attention on the MQAR [24] and LRA [3] benchmarks and train the three attention-based methods on WikiText-103. Inspired by S6, we define normalized attention as the attention function
@@
where $W_{\eta}\in\mathbb{R}^{1\times d}$ is an additional learnt parameter. In Appendix H we discuss two alternatives to (22). The MQAR results are shown in Figure 2 and the average accuracy on the LRA and the training perplexity on WikiText-103 in Table 1. The MQAR results suggest that proper normalization, i.e., using normalization (22), improves the performance of linear attention schemes. This is further supported by the performance of S6 and SSD on the MQAR benchmark, since these two methods also employ input-dependent normalization. Additionally, normalized attention closes part of the gap to softmax attention on the WikiText-103 dataset (Table 1). However on LRA, SSM models (S6) still achieve considerably higher performance than attention-based models. While normalized attention outperforms linear and softmax attention on the LRA, it performs significantly worse than S6. This result suggests that while the S6 inspired normalization helps to improve performance, the remaining performance gap is possibly explained by the recurrent normalization strategy employed by selective SSM models. Overall these results warrant further research into normalization strategies for attention-based models to explain the performance difference to SSMs. The complete experimental results on the MQAR and LRA benchmarks are detailed in Appendices L and M, respectively.
@@
Comparing RNNs and S6, it is immediate to observe several similarities. In particular, as shown in Appendix I, the state transition matrix $\Lambda_{i}$ in S6 (19) can be rewritten (assuming $A=a\cdot\mathbb{I}_{nd}$) as
@@
Notice that when no state expansion is considered, i.e., $n=1$ and $\mathbb{I}_{n}=1$, this expression almost coincides with the qLSTM state transition (20a), with the only difference that (I) it uses the reversed sigmoid function instead of a sigmoid for the forget gate, and (II) there is an additional learnt parameter $a$ in the exponent. Inspired by the subtle difference in the state transition, we compare the original qLSTM state transition (8) and the S6-inspired state transition 23 on the MQAR benchmark. The performance of both models is shown in Figure 3. We note that the reversed sigmoid state transition outperforms the original state transition on all three benchmark tasks, i.e., the performance of qLSTMs can be improved by insights from S6. Considering state expansion ($n>1$) for RNNs, the recent xLSTM paper [12] shows that state expanded LSTMs [^3] The presented state expanded LSTM versions, cannot be directly translated into the DSF framework, since the gates not only depend on the inputs $u_{i}$ but also on past outputs $y_{i-1}$. However, the used state expansion is essentially $n=d$, hence leading to a DSF system of size $d^{2}$. can yield similar performance to S6. This aligns with Lemma 2 and further highlights the importance of state expansion for expressivity. In qLSTM and RG-LRU, state expansion can be easily incorporated by changing the dimensions of the projections $W_{f},\,W_{i},\,W_{o}$, where the $\odot$ operation in RG-LRU would be replaced by blockwise operations $\otimes$. Finally, the most apparent difference between the two RNN variants – qLSTM and RG-LRU – and S6 is the parameter coupling in $\Lambda_{i}$ and $B_{i}$. While qLSTM does not use a coupling, the couplings in RG-LRU and S6 are performed with different nonlinearities, which is discussed in more detail in [10, Appendix A].
Figure 3: Comparison of qLSTM (8) and a qLSTM variant where the original state transition $\Lambda_{i}$ is replaced by (23).
@@ 5 Related Work
State-space models emerged from the S4 architecture by Gu et al. [25], who developed a new theoretically principled approach to sequence modeling rooted in polynomial approximation theory [26]. The result is a transformer-like architecture [2], where attention is replaced by a linear recurrent neural network with special reparametrization. The design of S4 got later simplified in [27, 4], achieving state-of-the-art performance on the long-range arena (LRA) [3] with a highly efficient recurrent mechanism leveraging convolutional views [28], or parallel scans [5, 6].
@@
The high efficiency of SSMs (linear processing) makes them particularly appealing when compared to softmax attention-based transformers, where both inference time and memory suffer quadratically from sequence length. The S4 architecture found first successful applications in reinforcement learning [29], vision [30], audio [31] as well as online learning [32]. Initial attempts in language modeling [33, 34], supported by theoretical investigations [35, 36] hint at some necessary architectural improvements to unlock the NLP domain. Leveraging in-context learning arguments, a few works [37, 38, 39] started incorporating input selectivity mechanisms [40] into SSMs. These efforts culminated in the Mamba architecture [7], which proposed a highly efficient and light (in terms of parameters) input selectivity strategy, with drastic improvements when comparing to earlier variants (H3 [33] and Hyena [41]) on text. This approach was also shown to be effective at byte level [42]. Beyond text, Mamba was recently applied to the vision domain [43, 44] – with outstanding results compared to ViTs [45] both in terms of performance and efficiency. Other applications include e.g. genetics [46], and point clouds [47]. Further, improvements on architectural aspects were proposed in [48, 8, 11].
@@
The design of Mamba is also strongly supported by theoretical evidence showing precisely its superior expressive power compared to S4 [49]. This boost in computational power is due to Mamba’s novel input selectivity mechanism resembling gating, which unlocks content-dependent reasoning [7, 40]. Interestingly, input selectivity brings SSMs closer to attention: in particular, Ali et al. [50] showed that the particular parametrization of Mamba can be linked to a non-normalized softmax operator. This finding is also supported by evidence from language theory – Mamba and Attention can solve a similar class of problems [51]. Beyond Ali et al. [50] the connection between linear attention and linear RNNs has been illustrated a few times in the literature [14, 52, 53, 23]. Connections between these architectures have also been carried out using tools from communication complexity in [54, 55]. Compared to these works and to Ali et al. [50], this paper offers a more careful comparison identifying some precise distinctions between SSMs, linear, and softmax attention – which play a nontrivial role in practice and can help bring to light interesting architectural improvements.